ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 30, 31, 35 Brodsky, Mikhail, University of California, Berkeley, USA A Physicist’s Definition of the Scwarz Function. The Schwarz function is a very interesting and helpful tool for solutions of several problems in complex analysis and partial differential equations. We present a definition of the function that arose in applied physical problems and gives a natural way for generalizations. Reference: P. Davis, The Schwarz Function and its Applications, Carus Mathematical Monographs 17, MAA, 1974.